OpenStudy (anonymous):

What is the fifteenth term of the geometric sequence -√5, √10, -2√5...? Show me how you got to your answer please!

OpenStudy (anonymous):

so the nth term is found by plugging that integer into n. but even though this sequence works. with only three terms, it can very well be the wrong one

OpenStudy (anonymous):

No, it just gives three terms. Then is says the possible answers are as follows: 1) -128√5 2) 128√10 3) -16384√5 4) 16384√10 I'm so lost. hahah

OpenStudy (anonymous):

\[(-1)^{n+1}\sqrt{5}\sqrt{2^n}\] starts at 0. 15th term: n = 14

OpenStudy (anonymous):

\[= (-1)^{14+1} \sqrt{5}\sqrt{2^{14}} = -\sqrt{5}\sqrt{16384}\] \[= -128\sqrt{5}\]

OpenStudy (anonymous):

so my choice was right. we can discuss a bit more how i chose it if u want, but it was really just trial and error and experience.

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

glad i could help :)